OpenStudy (anonymous):

The human resource department at a certain company wants to conduct a survey regarding worker benefits. The department has an alphabetical list of all 6606 employees at the company and wants to conduct a systematic sample of size 60 what is k?

OpenStudy (dumbcow):

the systematic sample forms an arithmetic series \[A = 1+k(n-1)\] let n = 60 A = 6606 solve for k, round down

OpenStudy (anonymous):

i got it now i think They ask me a second question tho. and i dont know how to answer it. Determine the individuals who will be administered the survey. Randomly select a number from 1 to k. suppose that we chose 17? Then the individuals in the survey will be( _,_,_)

OpenStudy (anonymous):

i got k=107

OpenStudy (dumbcow):

so if there are 6606 employees and k=17 the sample size will be 389 ---> 388*17 +1 = 6597